# Uncertainty-Aware Remaining Lifespan Prediction from Images

### v2

@@ Abstract
Predicting mortality-related outcomes from images offers the prospect of accessible, noninvasive, and scalable health screening. We present a method that leverages pretrained vision transformer foundation models to estimate remaining lifespan from facial and whole-body images, alongside robust uncertainty quantification. We show that predictive uncertainty varies systematically with the true remaining lifespan, and that this uncertainty can be effectively modeled by learning a Gaussian distribution for each sample. Our approach achieves state-of-the-art mean absolute error (MAE) of 7.48 years on an established dataset, and further improves to 4.79 and 5.07 years MAE on two new, higher-quality datasets curated and published in this work. Importantly, our models provide well-calibrated uncertainty estimates, as demonstrated by a bucketed expected calibration error of 0.62 years. While not intended for clinical deployment, these results highlight the potential of extracting medically relevant signals from images. We make all code and datasets available to facilitate further research.
@@ 1 Introduction
In this work, we propose a significantly improved approach to remaining lifespan prediction from images by combining powerful pre-trained vision transformers with a regression head that models prediction uncertainty as a Gaussian distribution. We show that such models produce well-calibrated uncertainty estimates, enabling better interpretation and downstream decision-making. Our contributions are as follows:
@@
  • We report state-of-the-art performance with an MAE of 4.79 years and an average calibration error of 0.62 years on the cleaned dataset.
@@ 2 Related Work
Epigenetic clocks and biological ageing. Estimating individual ageing rates from molecular data has been a major focus since Horvath’s seminal pan-tissue DNA-methylation clock [3]. Second-generation clocks such as DNAm PhenoAge [4] and DNAm GrimAge [5] improve correlations with disease burden and all-cause mortality, and are now widely used in geroscience studies. Crucially, however, these clocks are optimized for hazard ratios or risk scores rather than for point estimates of the time remaining to an individual, which limits their interpretability for person-centric counseling.
Image-based age and risk prediction. Computer-vision research on facial images initially targeted chronological age regression, achieving mean absolute errors below 2.5 years on the MORPH dataset [1, 2]. Fekrazad [6] collected a corpus of 24,000 Wikipedia portraits and trained convolutional models to regress remaining lifespan, reporting an MAE of 8.3 years. More recently, Bontempi et al. introduced FaceAge, a deep model that estimates biological age from a single headshot and enhances survival prediction in oncology cohorts [7]. While FaceAge improves survival hazard modeling, it does not produce calibrated estimates of absolute years of life remaining.
Other imaging modalities. Beyond faces, retinal fundus photographs [8], optical-coherence-tomography volumes [9], and chest radiographs [10, 11] have also been used to forecast longevity-related outcomes. Most of these pipelines feed image embeddings from a CNN (or Vision Transformer) into a Cox proportional-hazards loss—the classical survival-analysis objective that learns a log-linear function producing a hazard ratio—or its deep-learning analogue DeepSurv, which replaces the linear layer with an MLP while maintaining the same partial-likelihood formulation. Both objectives optimize the relative risk of death over time rather than predicting a concrete number of years left. Consequently, their outputs are survival curves or risk scores, underscoring that hazard modeling addresses a related, but distinct, question from the remaining-lifespan regression tackled in the present work.
Uncertainty-aware prediction. Reliable confidence estimates are indispensable for clinical deployment. Bayesian neural networks [12] and deep ensembles [13] have been proposed, but calibration for image-based lifespan prediction remains under-explored. Our Gaussian mean–variance head follows the mean-variance estimation paradigm [14, 15] and is, to our knowledge, the first to quantify uncertainty in image-based lifespan regression.
@@ 3 Dataset
  • Resolution and Color: The facial image must have a minimum resolution of 200$\times$200 pixels and contain color information.
@@ 5.1 Implementation Details
All training and evaluation runs were performed with a manually set PyTorch random seed (set to 1) to prevent random variations in results. We performed a standard 80/20 train-test split, resulting in 4,538 training and 1,134 test data points, respectively.
@@
We applied a standard normalization procedure to achieve approximately a mean and standard deviation of one for our target remaining lifespan data, and used an image normalization procedure similar to that used during backbone pre-training. For facial images, we resize to 224$\times$224 pixels according to DINOv2 pre-training. For whole-body images, we increase this to 1022$\times$1022, as we expect relevant details to be lost otherwise.
@@ 5.4 Results
Table 1: Test set results on different datasets and baselines.
@@
Ours (Faces Only) 4.79 0.62 -0.273
Ours (Whole Images) 5.07 0.85 -0.171
Ours (Legacy Dataset) 7.48 1.02 -0.388
Chronological Age Baseline (Faces Only) 7.80
Chronological Age Baseline (Legacy Dataset) 8.62
@@
Figure 4 provides a detailed analysis of model performance across different remaining lifespan ranges on the Faces Only dataset. The comparison between binned true and predicted errors demonstrates that our uncertainty estimates are well-calibrated across most buckets.
@@ 6 Discussion
Our method achieves state-of-the-art performance on remaining lifespan prediction, with an MAE of 7.48 years on the Legacy Dataset and substantially better results (4.79 and 5.07 years) on our newly curated datasets. However, this improvement is only partly attributable to higher data quality and improved modeling. The new datasets have lower average remaining lifespans, making the prediction task inherently easier—a side effect of filtering for high-quality images, which are more common for recently deceased individuals. Further research should prioritize curating more diverse datasets, for example from newspaper archives, to avoid this bias.
@@
The prediction error varies greatly across different remaining lifespan ranges, as expected due to the inherent aleatoric uncertainty of the task. Figure 4 shows that MAE, as a single-number metric, can be misleading when judging the reliability of the network for individual cases. In contrast, the uncertainty estimate produced by the network is well calibrated, with an average bucketed expected calibration error of just 0.62 years for the Facial Dataset.
@@
We consistently observe slightly worse results on the Whole Images Dataset. While most relevant information is expected to be contained in the face, we hypothesized that whole images might provide additional cues (e.g., weight, muscle mass) known to influence longevity, which can only be partially inferred from facial features. We thus conclude that the slight performance gap is inherent to the current approach. Since the DINOv2 backbone compresses the entire image into a single vector of size $d_{emb}=1536$, we expect it to contain less information about the face when processing whole images. If most relevant information is indeed contained in the face, this would explain the performance gap. However, even when we fine-tuned the backbone using an L1 loss (as described in Section 4.1), the performance gap persisted. We specifically provided higher resolution images to preserve facial information and allow the vision transformer backbone to focus on the most relevant parts of the image. We conclude that part of the performance loss is likely due to the pretraining of the DINOv2 model, which was trained on images of 224$\times$224 resolution [16], and does not generalize to larger image resolutions without some loss of performance.
@@ 7 Conclusion
We demonstrate that remaining lifespan prediction from facial and whole-body images is feasible and effective using pretrained vision foundation models. We further show that predictive uncertainty strongly depends on the remaining lifespan magnitude, and that this uncertainty can be reliably estimated, achieving a bucketed expected calibration error of 0.62 years. Our approach outperforms previous methods on the Legacy Dataset, and we curate and publish two new higher-quality datasets, achieving MAEs of 4.79 and 5.07 years, respectively. We hope that our research encourages further investigation into the extraction of medical information from images.

### v3

@@ Abstract
Predicting mortality-related outcomes from images offers the prospect of accessible, noninvasive, and scalable health screening. We present a method that leverages pretrained vision transformer foundation models to estimate remaining lifespan from facial and whole-body images, alongside robust uncertainty quantification. We show that predictive uncertainty varies systematically with the true remaining lifespan, and that this uncertainty can be effectively modeled by learning a Gaussian distribution for each sample. Our approach achieves state-of-the-art mean absolute error (MAE) of 7.41 years on an established dataset, and further achieves 4.91 and 4.99 years MAE on two new, higher-quality datasets curated and published in this work. Importantly, our models provide calibrated uncertainty estimates, as demonstrated by a bucketed expected calibration error of 0.82 years on the Faces Dataset. While not intended for clinical deployment, these results highlight the potential of extracting medically relevant signals from images. We make all code and datasets available to facilitate further research.
@@ 1 Introduction
In this work, we propose a significantly improved approach to remaining lifespan prediction from images by combining powerful pre-trained vision transformers with a regression head that models prediction uncertainty as a Gaussian distribution. We show that such models produce well-calibrated uncertainty estimates, enabling better interpretation and downstream decision-making. We evaluate all models using 5-fold cross-validation and report mean performance across folds. Our contributions are as follows:
@@
  • We report state-of-the-art performance with a mean MAE of 4.91 years and a mean calibration error of 0.82 years on the Faces Dataset.
@@ 2 Related Work
Epigenetic clocks and biological ageing. Estimating individual ageing rates from molecular data has been a major focus since Horvath’s seminal pan‑tissue DNA‑methylation clock [3]. Second‑generation clocks such as DNAmăPhenoAge [4] and DNAmăGrimAge [5] improve correlations with disease burden and all‑cause mortality, and are now widely used in geroscience studies. Crucially, however, these clocks are optimized for hazard ratios or risk scores rather than for point estimates of the time remaining to an individual, which limits their interpretability for person‑centric counseling.
Image‑based age and risk prediction. Computer‑vision research on facial images initially targeted chronological age regression, achieving mean absolute errors below 2.5 years on the MORPH dataset [1, 2]. Fekrazad [6] collected a corpus of 24,000 Wikipedia portraits and trained convolutional models to regress remaining lifespan, reporting an MAE of 8.3 years. More recently, Bontempi etăal. introduced FaceAge, a deep model that estimates biological age from a single headshot and enhances survival prediction in oncology cohorts [7]. While FaceAge improves survival hazard modeling, it does not produce calibrated estimates of absolute years of life remaining.
Other imaging modalities. Beyond faces, retinal fundus photographs [8], optical-coherence-tomography volumes [9], and chest radiographs [10, 11] have also been used to forecast longevity‐related outcomes. Most of these pipelines feed image embeddings from a CNN (or Vision Transformer) into a Cox proportional-hazards loss—the classical survival-analysis objective that learns a log-linear function producing a hazard ratio—or its deep-learning analogue DeepSurv, which replaces the linear layer with an MLP while maintaining the same partial-likelihood formulation. Both objectives optimize the relative risk of death over time rather than predicting a concrete number of years left. Consequently, their outputs are survival curves or risk scores, underscoring that hazard modeling addresses a related, but distinct, question from the remaining-lifespan regression tackled in the present work.
Uncertainty‑aware prediction. Reliable confidence estimates are indispensable for clinical deployment. Bayesian neural networks [12] and deep ensembles [13] have been proposed, but calibration for image-based lifespan prediction remains under‑explored. Our Gaussian mean–variance head follows the mean‑variance estimation paradigm [14, 15] and is, to our knowledge, the first to quantify uncertainty in image‑based lifespan regression.
@@ 3 Dataset
  • Resolution and Color: The facial image must have a minimum resolution of $\times200$ pixels and contain color information.
@@ 5.1 Implementation Details
All training and evaluation runs were performed with a manually set PyTorch random seed (set to 1) to prevent random variations in results. We evaluate using 5-fold cross-validation (KFold with shuffling, random state 1), reporting the mean test performance and standard error across folds.
@@
We applied a standard normalization procedure to achieve approximately a mean and standard deviation of one for our target remaining lifespan data, and used an image normalization procedure similar to that used during backbone pre-training. For facial images, we resize to $\times224$ pixels according to DINOv2 pre-training. For whole-body images, we increase this to $\times1022$, as we expect relevant details to be lost otherwise.
@@ 5.4 Results
Table 1: Mean test performance (5-fold cross-validation) and standard error on different datasets and baselines.
@@
Ours (Faces Only) 4.91 $\pm$ 0.2 0.82 $\pm$ 0.23 -0.216 $\pm$ 0.05
Ours (Whole Images) 4.99 $\pm$ 0.22 0.96 $\pm$ 0.26 -0.166 $\pm$ 0.07
Ours (Legacy Dataset) 7.41 $\pm$ 0.07 0.93 $\pm$ 0.08 -0.391 $\pm$ 0.02
CA Baseline (Faces Only) 7.80
CA Baseline (Legacy Dataset) 8.62
@@
Figure 4 provides a detailed analysis of model performance across different remaining lifespan ranges on the Faces Only dataset for the first fold. The comparison between binned true and predicted errors demonstrates that our uncertainty estimates are well-calibrated across most buckets.
@@ 6 Discussion
Our method achieves state-of-the-art performance on remaining lifespan prediction, with a mean MAE of 7.41 years on the Legacy Dataset and substantially better mean results (4.91 and 4.99 years) on our newly curated Faces and Whole Images datasets, respectively (all results via 5-fold cross-validation). However, this improvement is only partly attributable to higher data quality and improved modeling. The new datasets have lower average remaining lifespans, making the prediction task inherently easier—a side effect of filtering for high-quality images, which are more common for recently deceased individuals. Further research should prioritize curating more diverse datasets, for example from newspaper archives, to avoid this bias.
@@
The prediction error varies greatly across different remaining lifespan ranges, as expected due to the inherent aleatoric uncertainty of the task. Figure 4 shows that MAE, as a single-number metric, can be misleading when judging the reliability of the network for individual cases. In contrast, the uncertainty estimate produced by the network is well calibrated on average, with a mean bucketed expected calibration error of 0.82 years on the Faces Dataset (95% CI [0.59, 1.05]).
@@
We consistently observe slightly worse results on the Whole Images Dataset (MAE 4.99 vs. 4.91 on Faces). While most relevant information is expected to be contained in the face, we hypothesized that whole images might provide additional cues (e.g., weight, muscle mass) known to influence longevity, which can only be partially inferred from facial features. We thus conclude that the slight performance gap is inherent to the current approach. Since the DINOv2 backbone compresses the entire image into a single vector of size $d_{emb}=1536$, we expect it to contain less information about the face when processing whole images. If most relevant information is indeed contained in the face, this would explain the performance gap. However, even when we fine-tuned the backbone using an L1 loss (as described in Section 4.1), the performance gap persisted. We specifically provided higher resolution images to preserve facial information and allow the vision transformer backbone to focus on the most relevant parts of the image. We conclude that part of the performance loss is likely due to the pretraining of the DINOv2 model, which was trained on images of $\times224$ resolution [16], and does not generalize to larger image resolutions without some loss of performance.
@@ 7 Conclusion
We demonstrate that remaining lifespan prediction from facial and whole-body images is feasible and effective using pretrained vision foundation models. We further show that predictive uncertainty strongly depends on the remaining lifespan magnitude, and that this uncertainty can be reliably estimated, achieving a mean bucketed expected calibration error of 0.82 years on the Faces Dataset. Our approach outperforms previous methods on the Legacy Dataset, and we curate and publish two new higher-quality datasets, achieving mean MAEs of 4.91 (Faces) and 4.99 (Whole Images) years, respectively, under 5-fold cross-validation. We hope that our research encourages further investigation into the extraction of medical information from images.